# VEAttack: Downstream-agnostic Vision Encoder Attack against Large Vision Language Models

### v1

@@ Appendix E Evaluation on image classification datasets
Since CLIP can also function as a classifier, we assess the impact of VEAttack across diverse classification datasets, including Tiny-ImageNet [15], CIFAR-10 [26], CIFAR-100 [26], FOOD101 [6], Flowers102 [40], DTD [12], EuroSAT [22], FGVC-Aircraft [37], Caltech-256 [20], StanfordCars [25], ImageNet [15] and SUN397 [60]. Table 8 presents the evaluation of VEAttack on image classification datasets using the CLIP-B/32 model. The results show that $\mathcal{L}_{cls}$, which optimizes the cross-entropy loss between classification and text features, achieves the most significant performance degradation of $97.8\%$, as its objective aligns closely with the classification task. Compared with $\mathcal{L}_{2}$ attack, which uses the $\ell_{2}$ loss of class token embedding as the attack objective, VEAttack achieves a more substantial reduction in mean accuracy as $93.0\%$, demonstrating its superior efficacy in disrupting classification performance. This improvement also highlights the advantage of perturbing image tokens over class tokens, underscoring its effectiveness over traditional methods.
@@
$\mathcal{L}_{cls}$ [39] $\epsilon=4/255$ 0.5 2.1 0.1 5.4 0.4 0.0 0.0 0.0 6.8 0.1 0.4 0.3 1.3
Appendix F Qualitative results of VEAttack
F.1 Qualitative results on image captioning
We visualize the outputs of clean samples, those attacked by the traditional white-box APGD method [13], and our proposed VEAttack, as shown in Figure 8. Both attack methods cause errors in the image captioning outputs of LVLMs. However, an interesting observation emerges: the traditional white-box attack, which targets output tokens as its objective, tends to generate errors that still revolve around the original sentence structure or subject, such as "in the water" in the first image, "running" in the second, and "spraying water" in the third. In contrast, since VEAttack operates in a downstream-agnostic context, it directly disrupts the semantic information in visual features of the vision encoder, leading to outputs that completely deviate from the clean captions, for example, "table, bear" in the first image, "food" in the second, and "game controller" in the third. While recent work [44] mitigates such issues through targeted attack to improve effectiveness, our non-targeted VEAttack approach successfully achieves a complete semantic deviation under non-targeted conditions.
F.2 Qualitative results on visual question answering
We visualize the outputs of LVLMs on the visual question answering task under clean, APGD, and VEAttack conditions, as shown in Figure 9. Notably, VEAttack, despite operating in a downstream-agnostic context without instruction-specific guidance, effectively disrupts visual features, leading to erroneous responses from the model in most cases. This highlights its robust attack capability. We also present a failure case in the last column, where VEAttack does not specifically alter the semantic information related to the brand monitor. Overall, VEAttack demonstrates exceptional attack effectiveness in a downstream-agnostic manner.

### v2

@@ Appendix I Transfer attack from larger specific vision encoders
Since the CLIP can also function as a classifier, we assess the impact of VEAttack across diverse classification datasets, including Tiny-ImageNet (Deng et al., 2009), CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009), FOOD101 (Bossard et al., 2014), Flowers102 (Nilsback and Zisserman, 2008), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), FGVC-Aircraft (Maji et al., 2013), Caltech-256 (Griffin et al., 2007), StanfordCars (Krause et al., 2013), ImageNet (Deng et al., 2009) and SUN397 (Xiao et al., 2010). Table 14 presents the evaluation of VEAttack on image classification datasets using the CLIP-B/32 model. The results show that $\mathcal{L}_{cls}$, which optimizes the cross-entropy loss between classification and text features, achieves the most significant performance degradation of $97.8\%$, as its objective aligns closely with the classification task. Compared with the $\mathcal{L}_{2}$ attack, which uses the $\ell_{2}$ loss of class token embedding as the attack objective, VEAttack achieves a more substantial reduction in mean accuracy of $93.0\%$, demonstrating its superior efficacy in disrupting classification performance. This improvement also highlights the advantage of perturbing patch tokens over class tokens, underscoring its effectiveness over traditional methods.
@@
$\mathcal{L}_{cls}$ Mao et al. (2023) $\epsilon=4/255$ 0.5 2.1 0.1 5.4 0.4 0.0 0.0 0.0 6.8 0.1 0.4 0.3 1.3
Appendix K Qualitative results of VEAttack
K.1 Comparison with perturbation of black box attacks
Fig. 10 shows the adversarial images generated by our VEAttack, and two targeted black-box attacks, AnyAttack (Zhang et al., 2024) and FOA-Attack (Jia et al., 2025). Visual inspection shows that VEAttack yields substantially higher imperceptibility. For example, in the second row, the adversarial results of AnyAttack and FOA‑Attack imprint recognizable fruit contours from the target image onto the green leaf, making the manipulation conspicuous, whereas the VEAttack image remains nearly indistinguishable from the clean input. To make this contrast explicit, we also visualize the perturbations. VEAttack achieves effective LVLM degradation with low‑magnitude, semantically unstructured noise, indicating that VEAttack could attain stronger perceptual stealth.
Figure 10: Comparison with perturbation of targeted black-box attacks AnyAttack (Zhang et al., 2024) and FOA-Attack (Jia et al., 2025).
K.2 More Comparisons of Adversarial Images
Figure 11: Visual comparison of adversarial examples generated by the black-box M-Attack (Li et al., 2025) and our gray-box VEAttack.
To demonstrate the superior imperceptibility of VEAttack, we conduct a detailed qualitative and quantitative comparison with the black-box M-Attack (Li et al., 2025), as shown in Figure 11. Our goal regarding imperceptibility is to achieve effective attacks using significantly smaller perturbations compared to black-box methods, which often require heavy noise to ensure transferability. Visually, M-Attack introduces conspicuous texture distortions. Notably, it tends to generate perturbations containing semantic information; for instance, in the fourth column of the first row, word-shaped artifacts are clearly visible in the sky. In contrast, VEAttack generates perturbations that are texture-agnostic and devoid of semantic artifacts, making them much harder for human observers to detect. Quantitatively, we evaluated 500 randomly sampled adversarial images. VEAttack achieves an average $L_{2}$ distance of 4.3883 between the adversarial and clean images, whereas M-Attack yields a much higher distance of 15.3403. Furthermore, we calculate the cosine similarity (CLIP Score) between the CLIP-L/14 features of the clean and adversarial images. VEAttack achieves a score of 0.4443, compared to 0.5493 for M-Attack. This indicates that VEAttack successfully pushes the visual features further away from the original semantics, while maintaining a much lower visual perturbation budget than the black-box attacks.
@@
K.3 Qualitative results on image captioning
We visualize the outputs of clean samples, those attacked by the traditional white-box APGD method (Croce and Hein, 2020), and our proposed VEAttack, as shown in Fig. 12. Both attack methods cause errors in the image captioning outputs of LVLMs. However, an interesting observation emerges: the traditional white-box attack, which targets output tokens as its objective, tends to generate errors that still revolve around the original sentence structure or subject, such as ”in the water” in the first image, ”running” in the second, and ”spraying water” in the third. In contrast, since VEAttack operates in a downstream-agnostic context, it directly disrupts the semantic information in visual features of the vision encoder, leading to outputs that completely deviate from the clean captions, for example, ”table, bear” in the first image, ”food” in the second, and ”game controller” in the third. While recent work (Schlarmann and Hein, 2023) mitigates such issues through targeted attacks to improve effectiveness, our non-targeted VEAttack approach successfully achieves a complete semantic deviation under non-targeted conditions.
K.4 Qualitative results on visual question answering
We visualize the outputs of LVLMs on the visual question answering task under clean, APGD, and VEAttack conditions, as shown in Fig. 13. Notably, VEAttack, despite operating in a downstream-agnostic context without instruction-specific guidance, effectively disrupts visual features, leading to erroneous responses from the model in most cases. This highlights its robust attack capability. We also present a failure case in the last column, where VEAttack does not specifically alter the semantic information related to the brand monitor. Overall, VEAttack demonstrates exceptional attack effectiveness in a downstream-agnostic manner.
Appendix L The use of Large Language Models
In accordance with the policies on LLM usage, we used a large language model as a language‑refinement assistant during manuscript preparation. The use of the LLM was confined strictly to editorial support and did not contribute to the conceptual or experimental aspects of this research. Specifically, the LLM was employed for:
  • Rephrase sentences and paragraphs to improve clarity, concision, and academic tone when describing our gray‑box, vision‑encoder‑only attack, its objectives, and empirical observations (e.g., the shift from class‑token to patch‑token perturbations and the cosine‑similarity objective)
  • Standardize terminology and notation across sections (e.g., downstream‑agnostic, untargeted, limited perturbation budget, and the symbols in Eqs. (4)–(7))
  • Smooth local transitions (e.g., from the redefined objective in Section 4.1 to the loss design in Section 4.2) and polish captions for clarity.
Our research ideas, technical methods, proofs, experiments, figures, and conclusions are entirely the work of the human authors. All LLM‑suggested text was reviewed line‑by‑line and edited by the authors. Any technical statements, equations, or quantitative claims appearing in the paper were verified against our experiments and, where appropriate, cross‑checked with the paper’s own tables and figures.